Grade 12Algebra

A car runs 16 km using 1 litre of petrol. How much distance will it cover using 2 ¾ litres of petrol.

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

From the question, it is given that,
The total number of distance travelled by a car in 1 liter of petrol = 16 km
Then,
Total quantity of petrol = 2 whole 3⁄4 liter
= 11/4 liters
Total number of distance travelled by car in 11/4 liters of petrol = (11/4) × 16
= 11 × 4
= 44 km
∴Total number of distance travelled by car in 11/4 liters of petrol is 44 km.

Thanks